### v1

@@ 2.6 Model Training Configuration
2.7 Cross-Validation and Generalization Assessment
@@ 2.7.1 Feature Scaling and Normalization
All features were standardized using user-specific z-score normalization following established protocols for cross-participant physiological models (bhatti2024clare). The scaler was fitted on the training data and applied consistently during real-time deployment to ensure feature distributions matched those seen during training. This approach addresses the critical challenge of cross-participant generalization in physiological signal classification, where baseline differences in heart rate, pupil size and skin conductance can significantly impact model performance.
@@ 3.1 Model Evaluation (pre-deployment)
Feature Importance Analysis: Permutation feature attribution analysis (fisher2019all) revealed that oculometric measures dominated model predictions, with all top 10 contributing features derived from eye-tracking data (Figure 5c). The most informative features comprised saccadic metrics (6 features: velocity, amplitude, frequency, and direction characteristics), fixation-related parameters (2 features: duration and spatial dispersion), pupil diameter variability, and blink frequency patterns.
@@ 3.2.1 Overall Classification Performance
The deployed model achieved 36.7% overall accuracy in predicting participant difficulty ratings during real-time operation, representing a substantial performance degradation from the 52.3% offline training accuracy. This performance gap highlights the inherent challenges of translating laboratory-trained models to dynamic, interactive environments where temporal context, user adaptation, and real-time processing constraints impact classification reliability. The confusion matrix (Figure 6a) revealed systematic bias toward "Too Easy" predictions, with the most common error being classification of "Just Right" trials incorrectly as "Too Easy" (30.3% of all trials). Class-specific precision demonstrated the anticipated training strategy effects: moderate precision for "Just Right" classifications (0.60), reflecting the model’s design emphasis on edge-case detection, while "Too Easy" (0.31) and "Too Difficult" (0.27) categories showed lower precision values.
@@ 3.4 Game Performance Analysis: Influences of Control Mechanism and Task Complexity
Measuring game performance in adaptive virtual reality tasks presents unique challenges, particularly when task difficulty varies dynamically based on participant ability. Here we use three complementary metrics of game performance to obtain a nuanced view of performance:
@@
A 2×2 repeated measures ANOVA was conducted with Control Type (Model vs Self) and Task Complexity (Single vs Dual) as within-subjects factors and raw game performance as well as difficulty-weighted game performance as dependent variables. The key results for this section are visualised in Figure 8. Raw game performance analysis revealed a significant Control × Task Complexity interaction (F(1,50) = 9.387, p = 0.003), with no significant main effects for Control Type (F(1,50) = 2.213, p = 0.143) or Task Complexity (F(1,50) = 3.076, p = 0.086). The interaction pattern demonstrated that model guidance resulted in equivalent raw game performance to self-control in single-task conditions (0.909 vs 0.903), while self-control demonstrated superior raw game performance in dual-task conditions (0.933 vs 0.897).
Critically, difficulty-weighted game performance analysis revealed a significant main effect for Task Complexity (F(1,50) = 4.045, p = 0.050), indicating that game performance differed between single and dual tasks when accounting for task difficulty, with no significant effects for Control Type (F(1,50) = 1.554, p = 0.218) or the Control × Task Complexity interaction (F(1,50) = 0.058, p = 0.811). This suggests that while the adaptive system maintained equivalent challenge calibration across control types, task complexity influenced game performance outcomes when weighted by difficulty.
@@
3.5 Evaluating the Real-time Model based on Adaptive Difficulty Control Effectiveness - Analysing the Final Difficulty Convergence
@@
The individual predictions might not be the most reliable way of evaluating the accuracy of the model, given that they rely on the participant and model agreeing about individual 60s-long level. That method of evaluation, although providing an absolute measure of accuracy, misses out on the general trend in difficulty adjustment. We argue that if the model converges to a similar difficulty level as the participant does by the end of each condition, then it is an indication that the overall cognitive load trend has been captured accurately. Conversely, if the model diverges from the participant-driven difficulty, it should be evidence for failure of the adaptive system. Therefore, we investigated the patterns of difficulty convergence between model and participant driven conditions.
The interaction between control type and task complexity proved crucial for understanding adaptive model effectiveness. A 2×2 repeated measures ANOVA was conducted with Control Type (Model vs Self) and Task Complexity (Single vs Dual) as within-subjects factors and the difficulty of the final level played in each condition as the dependent variable. The results revealed no significant main effect of task complexity ($t(50)=1.62$, $p=0.112$), and the main effect of control type approached but did not reach significance (t(50)=1.85t(50)=-1.85, $p=0.070$). Although the omnibus interaction was not statistically significant (t(50)=1.79t(50)=-1.79, $p=0.080$), planned simple main effects analysis was conducted to test the specific hypothesis regarding model behavior under load. In single-task conditions, model-based and self-assessment control achieved virtually equivalent final difficulty levels (Model: $1.76\pm 0.55$, Self: $1.79\pm 0.38$, t=0.37t=-0.37, $p=0.714$), indicating comparable performance when cognitive load was low. However, in dual-task conditions, model-based control assigned significantly harder final difficulties than participants selected for themselves (Model: $1.73\pm 0.60$, Self: $1.98\pm 0.53$, t=2.31t=-2.31, $p=0.025$, $d=0.33$). After applying a Bonferroni correction for two comparisons, this difference remained significant (corrected $p=0.05$), see Figure 9. This divergence, along with no difference in weighted game performance across dual tasks, indicates that participants are more conservative with their assessments, playing lower dual task levels than when guided by the model but achieving similar weighted game performance and equivalent arithmetic error rates. This suggests that despite the model pushing the participants to play at harder levels in dual task designs, it does not come at a cost in difficulty-weighted performance.
@@ 3.7 Subjective Workload Assessment with NASA TLX
While real-time cognitive workload evaluation based on categorical difficulty labels (Too Easy, Just Right, Too Difficult) provides insight into moment-to-moment task adaptation, it may not fully capture the broader experiential and affective dimensions of user engagement. To complement this model-driven assessment, we employed the NASA Task Load Index (TLX) at the end of each experimental condition. This validated instrument assesses participants’ subjective workload across six dimensions: mental demand, physical demand, temporal demand, performance perception, effort, and frustration. A comprehensive within-participant analysis was conducted to examine workload differences between control types (Model vs Self) and task complexity (Dual vs Single).
3.7.1 Task Complexity Effects in NASA TLX ratings
The addition of the arithmetic secondary task produced substantial and consistent increases in subjective workload across multiple dimensions. Dual task conditions resulted in significantly higher workload ratings compared to single-task conditions for mental demand ($\Delta=4.25$, $t=7.725$, $p<0.001$, $d=1.41$, $\text{SD}_{\text{Dual}}=3.7$, $\text{SD}_{\text{Single}}=4.3$), effort ($\Delta=2.55$, $t=4.359$, $p<0.001$, $d=0.78$, $\text{SD}_{\text{Dual}}=3.8$, $\text{SD}_{\text{Single}}=5.1$), and frustration ($\Delta=2.45$, $t=4.013$, $p<0.001$, $d=0.69$, $\text{SD}_{\text{Dual}}=5.0$, $\text{SD}_{\text{Single}}=4.8$). Physical demand ($\Delta=1.05$, $t=2.619$, $p=0.012$, $d=0.46$, $\text{SD}_{\text{Dual}}=4.8$, $\text{SD}_{\text{Single}}=5.2$) and temporal demand ($\Delta=1.20$, $t=1.807$, $p=0.077$, $d=0.36$, $\text{SD}_{\text{Dual}}=3.7$, $\text{SD}_{\text{Single}}=4.8$) also showed increases, with physical demand reaching significance but temporal demand showing a non-significant trend. Total workload scores were significantly higher in dual task conditions ($\Delta=11.70$, $t=4.531$, $p<0.001$, $d=0.87$, $\text{SD}_{\text{Dual}}=17.2$, $\text{SD}_{\text{Single}}=20.9$), demonstrating the substantial cognitive cost imposed by concurrent arithmetic processing.
3.7.2 Control Type Effects in NASA TLX ratings
No significant differences emerged between Model and Self control conditions across most TLX dimensions. Mental demand showed a marginal trend ($\Delta=0.90$, $t=1.824$, $p=0.074$, $d=0.26$, $\text{SD}_{\text{Model}}=3.9$, $\text{SD}_{\text{Self}}=4.1$), with Model control slightly higher than Self control, but this difference failed to reach statistical significance. Similarly, performance perception showed a non-significant trend favoring Self control ($\Delta=0.60$, $t=1.287$, $p=0.204$, $d=0.18$, $\text{SD}_{\text{Model}}=3.8$, $\text{SD}_{\text{Self}}=3.2$). Total workload scores were comparable between control types (Model: $60.7\pm 14.9$, Self: $57.7\pm 19.5$, $t=1.563$, $p=0.124$, $d=0.22$), indicating that participants experienced similar subjective burden regardless of whether difficulty was adjusted through model predictions or self-assessment.
@@ 3.8 Learning and Fatigue Effects
To assess participant adjustment to the task over time, we analyzed learning and fatigue patterns by comparing participants’ game performance and NASA-TLX results between their first and second exposure to each task type (single-task vs dual-task conditions). The experiment design ensured consistent sequencing across participants. Single task blocks were always in positions 1 and 3, while dual task blocks appeared in positions 2 and 4. This fixed ordering allowed for direct within-subject comparisons, controlling for time-on-task and interval effects across all participants. Repeated-measures ANOVA was conducted separately for each task type with single-task results presented in Table LABEL:tab:single_learning and dual-task results presented in Table LABEL:tab:dual_learning.
@@
3.8.1 Game Performance Learning Effects
Weighted game performance showed significant improvement from first to second exposure for both task types, with larger effects for single-task conditions. Single-task weighted game performance increased by 0.224 points (p < 0.001, Cohen’s d = 0.866), while dual-task weighted game performance improved by 0.146 points (p < 0.001, d = 0.532). Raw game performance improvements were more modest, with dual-task conditions showing a significant but small increase of 0.024 points (p = 0.016, d = 0.348), while single-task raw game performance changes were not significant (p = 0.075). Figure 12a demonstrates these weighted game performance learning effects, revealing that participants achieved better game performance while maintaining or progressing to more challenging difficulty levels on their second exposure to each task type.
3.8.2 Difficulty Progression and Perceived Challenge Over Time
Participants achieved significantly harder final difficulty levels on their second single-task exposure, with final difficulty decreasing by 0.162 points (p < 0.001, d = 0.530), as shown in Figure 12c, indicating learning effects. No significant change was observed in dual-task final difficulty (p = 0.108).

### v2

@@ 2.6 Model Training Configuration
2.7 Cross-Validation, Generalisability & Interpretability
@@ 2.7.1 Feature Scaling and Normalization
All features were standardized using a global z-score normalization strategy, where scaling parameters were derived from the training set. These fixed parameters were applied directly to the real-time data stream during deployment, treating the training distribution as representative of the target population. This approach was selected to facilitate a zero-calibration framework, enabling the system to generate immediate, one-shot predictions for novel users without the requirement for obtaining pre-session baselines or participant-specific adaptation.
2.7.2 Feature importance analysis
To interpret the key decision markers behind our model, we used SHAP (SHapley Additive exPlanations) (lundberg2017unified), which is a versatile approach used to interpret the results of machine learning models. It calculates the contribution of each feature to the prediction of a specific instance, regardless of the model used. This method leverages Shapley values from cooperative game theory to distribute the overall value generated by a group of players fairly among individual players. By assigning SHAP values to each feature, one can effectively rank the features based on their importance in the prediction process (lundberg2017unified). As opposed to traditional explainability approaches which estimate feature importance by shuffling one feature at a time, SHAP is relatively resilient to correlated variables.
@@ 3.1 Model Evaluation (pre-deployment)
Feature Importance Analysis: The SHAP analysis of feature importance revealed that oculometric measures dominated model predictions, with the top 7 contributing features derived exclusively from eye-tracking data (Figure 5c). The most informative features were primarily saccadic metrics (specifically mean acceleration, peak velocity, and mean amplitude) and fixation-related parameters (standard deviation of duration and count), alongside pupil diameter variability. Physiological measures played a secondary role, with short-term Heart Rate Variability (RMSSD) appearing as the eighth most important feature, while blink frequency patterns were notably absent from the top contributions.
@@ 3.2.1 Overall Classification Performance
The deployed model achieved 36.7% overall accuracy in predicting participant difficulty ratings during real-time operation, representing a substantial performance degradation from the 52.3% offline training accuracy. This performance gap highlights the inherent challenges of translating laboratory-trained models to dynamic, interactive environments where temporal context, user adaptation, and real-time processing constraints impact classification reliability. The confusion matrix (Figure 6a) revealed systematic bias toward "Too Easy" predictions, with the most common error being classification of Just Right trials incorrectly as Too Easy (30.3% of all trials). Class-specific precision demonstrated the anticipated training strategy effects: moderate precision for Just Right classifications (0.60), reflecting the model’s design emphasis on edge-case detection, while Too Easy (0.31) and Too Difficult (0.27) categories showed lower precision values.
@@ 3.4 Game Performance Analysis: Influences of Control Mechanism and Task Complexity
Measuring game performance in adaptive virtual reality tasks is a nuanced problem, particularly when task difficulty varies dynamically based on participant ability. Here we use three complementary metrics of game performance to obtain a robust view of performance:
@@
A series of 2×2 repeated measures ANOVAs was conducted with Control Type (Model vs Self) and Task Complexity (Single vs Dual) as within-subjects factors and raw game performance as well as difficulty-weighted game performance as dependent variables. A paired t-test was used to determine whether there was a difference in arithmetic errors in the dual conditions depending on Control Type. Prior to analysis, assumption checks were performed for the primary dependent variables. Outlier detection utilizing the Interquartile Range (IQR) method identified a limited number of outliers across conditions. Shapiro-Wilk tests indicated that raw game performance data violated the assumption of normality ($p<.001$) due to ceiling effect expected in levels that were too easy and did not sufficiently challenge participants. To ensure statistical rigor, we conducted the primary analysis using standard repeated measures ANOVA and validated significant findings using a robustness check with arcsine square-root transformed data. The key results for this section are visualised in Figure 8. Analysis of raw game performance revealed a significant Control Type $\times$ Task Complexity interaction ($F(1,50)=7.16,p=.008$). While performance was statistically equivalent in single-task conditions ($M_{Model}=0.91$ vs. $M_{Self}=0.90$), participants in the self-control condition achieved significantly higher raw accuracy in dual-task conditions ($M_{Self}=0.93$) compared to the model-control condition ($M_{Model}=0.90$). This interaction effect was confirmed by the robustness check on transformed data ($F(1,50)=11.44,p=.001$), indicating that the divergence in strategies is a robust phenomenon and not an artifact of the data distribution.
Critically, difficulty-weighted game performance analysis revealed a significant main effect for Task Complexity ($F(1,50)=4.045,p=0.050$), indicating that game performance differed between single and dual tasks when accounting for task difficulty, with no significant effects for Control Type ($F(1,50)=1.55,p=.219$) or the Control × Task Complexity interaction ($F(1,50)=4.04,p=.050$). This suggests that while the adaptive system maintained equivalent challenge calibration across control types, task complexity influenced game performance outcomes when weighted by difficulty.
@@
3.5 Adaptive Control Effectiveness: Difficulty Trends and Convergence
@@
The individual predictions might not be the most reliable way of evaluating the accuracy of the model, given that they rely on the participant and model agreeing about individual 60s-long level. That method of evaluation, although providing an absolute measure of accuracy, misses out on the general trend in difficulty adjustment. We argue that if the model converges to a similar difficulty level as the participant does by the end of each condition, then it is an indication that the overall cognitive load trend has been captured accurately. Conversely, if the model diverges from the participant-driven difficulty, it should be evidence for failure of the adaptive system. Therefore, we investigated both the temporal dynamics of difficulty adjustment and the final convergence points between conditions.
To capture the continuous nature of this adaptation process, we first analyzed the slope of difficulty changes over time using a Linear Mixed-Effects Model (LMM). The model included fixed effects for Time (Level Sequence), Control Type (Model vs Self), and Task Complexity (Single vs Dual), along with random intercepts and slopes for Time grouped by Participant. Residual diagnostics indicated a minor deviation from normality ($W=0.953$). This deviation was attributed to the discrete, step-wise nature of the difficulty levels. Given the large number of observations ($N=1,632$) and the balanced study design, the model was deemed robust to these minor violations without the need for transformation. We followed that by a $2\times 2$ repeated measures ANOVA was conducted on the final level played, with Control Type and Task Complexity as factors. Prior to analysis, the distribution of final difficulty levels was inspected. Shapiro-Wilk tests indicated deviations from normality in all conditions ($W$ range: 0.830.950.83-0.95, $p<.05$). Visual inspection of Q-Q plots revealed that these deviations were driven by the discrete nature of the difficulty levels (step-wise increments) rather than extreme outliers. Given the balanced design and sample size ($N=51$), the repeated measures ANOVA was deemed robust to these violations.
The analysis revealed a significant main effect of Time ($\beta=-0.122,SE=0.009,p<.001$), confirming that objective difficulty generally increased (numerical value decreased) as the session progressed across all conditions. Crucially, a marginally significant three-way interaction was observed between Time, Control Type, and Task Complexity ($\beta=0.024,SE=0.013,p=.073$). As visualized in Figure 9c and 9d, the difficulty trajectories diverge under load. While the participants’ adaptation slope was aligned with the model’s in the single task condition (Figure 9c), it diverged in the dual-task condition (Figure 9d, red line), increasing difficulty at a slower rate than in the model-controlled condition.
This divergence in adaptation rates accumulated over the session, culminating in a crucial interaction at the final difficulty level. The results of the $2\times 2$ repeated measures ANOVA revealed no significant main effect of task complexity ($t(50)=1.62$, $p=0.112$), and the main effect of control type approached but did not reach significance (t(50)=1.85t(50)=-1.85, $p=0.070$). Although the omnibus interaction was not statistically significant (t(50)=1.79t(50)=-1.79, $p=0.080$), planned simple main effects analysis confirmed the trend observed in the slope analysis.
In single-task conditions, model-based and self-assessment control achieved virtually equivalent final difficulty levels (Model: $1.76\pm 0.55$, Self: $1.79\pm 0.38$, t=0.37t=-0.37, $p=0.714$), indicating comparable performance when cognitive load was low (see Figure 9a). However, in dual-task conditions, model-based control assigned significantly harder final difficulties than participants selected for themselves (Model: $1.73\pm 0.60$, Self: $1.98\pm 0.53$, t=2.31t=-2.31, $p=0.025$, $d=0.33$). After applying a Bonferroni correction for two comparisons, this difference remained significant (corrected $p=0.05$).
The divergence in final difficulty played, along with no difference in weighted game performance across dual tasks, indicates that participants are more conservative with their assessments. They adopted a flatter adaptation slope (Figure 9d) and ended at easier levels than the model determined they were capable of handling but ended up achieving similar weighted game performance and equivalent arithmetic error rates. This suggests that despite the model pushing the participants to play at harder levels in dual task designs, it does not come at a cost in difficulty-weighted performance.
@@ 3.7 Subjective Workload Assessment with NASA TLX
While real-time cognitive workload evaluation based on categorical difficulty labels (Too Easy, Just Right, Too Difficult) provides insight into moment-to-moment task adaptation, it may not fully capture the broader experiential and affective dimensions of user engagement. To complement this model-driven assessment, we employed the NASA Task Load Index (TLX) at the end of each experimental condition.
To examine the effects of the adaptive system and the secondary task on subjective experience, a $2\times 2$ repeated measures ANOVA was conducted on the weighted TLX scores, with Control Type (Model vs. Self) and Task Complexity (Single vs. Dual) as within-subject factors. Shapiro-Wilk tests confirmed that NASA-TLX workload scores satisfied the assumption of normality across all four experimental conditions ($W>0.96$, $p>.19$), validating the use of parametric analysis without correction.
3.7.1 Task Complexity
The analysis revealed a robust main effect of Task Complexity across nearly all workload dimensions, confirming the efficacy of the arithmetic task in inducing cognitive load. For the overall Total Workload score, participants reported significantly higher burden in the dual-task condition compared to the single-task condition ($F(1,50)=38.48,p<.001$).
Breakdown by subscale indicated that this increase was driven primarily by Mental Demand ($F(1,50)=101.03,p<.001$) and Effort ($F(1,50)=31.30,p<.001$), reflecting the high cognitive cost of concurrent arithmetic processing. Frustration levels were also significantly higher in the dual-task condition ($F(1,50)=24.28,p<.001$), suggesting that the divided attention requirement negatively impacted affective state. Physical Demand ($F(1,50)=10.87,p=.002$) and Temporal Demand ($F(1,50)=6.45,p=.014$) showed smaller but statistically significant increases. notably, the Performance subscale did not differ significantly between task conditions ($F(1,50)=2.40,p=.127$), indicating that participants felt their level of success was comparable regardless of the added complexity.
3.7.2 Control Type and Interaction Effects
In contrast to the strong effects of task complexity, the main effect of Control Type was not statistically significant for Total Workload ($F(1,50)=1.87,p=.177$), nor for any individual subscale (all $p>.08$). This indicates that participants did not perceive the Model-controlled adaptive difficulty as significantly more demanding or intrusive than their own self-regulated adjustments.
Crucially, the interaction between Control Type and Task Complexity was not significant for Total Workload ($F(1,50)=0.27,p=.608$) or any subscale. This suggests that the increase in workload observed during the dual-task condition was consistent across both control methods. The adaptive model did not exacerbate the workload of the secondary task, nor did it mitigate it significantly more than users could achieve on their own.
@@ 3.8 Learning and Fatigue Effects
To assess participant adjustment to the task over time, we analysed learning and fatigue patterns by comparing participants’ game performance and NASA-TLX results between their first and second exposure to each task type (single-task vs dual-task conditions). The experiment design ensured consistent sequencing across participants. Single task blocks were always in positions 1 and 3, while dual task blocks appeared in positions 2 and 4. This fixed ordering allowed for direct within-subject comparisons, controlling for time-on-task and interval effects across all participants. Repeated-measures ANOVA was conducted separately for each task type with single-task results presented in Table LABEL:tab:single_learning and dual-task results presented in Table LABEL:tab:dual_learning.
@@
3.8.1 Performance Dynamics and Skill Acquisition
To assess the sustainability of performance and the presence of learning effects over the course of the session, we analysed the trajectory of difficulty-weighted performance. A Linear Mixed-Effects Model (LMM) was fitted with Level Sequence (Time), Control Type, and Task Complexity as fixed effects, including random intercepts and slopes for participants. Prior to analysis, residual diagnostics were performed to validate the model. The Shapiro-Wilk test on model residuals indicated an excellent fit to the normal distribution ($W=0.997$), and visual inspection of Q-Q plots confirmed that residuals closely followed the theoretical diagonal. A scatterplot of residuals versus fitted values displayed a random dispersion of points, satisfying the homoscedasticity assumption required for linear mixed-effects modeling.
The analysis revealed a robust and statistically significant main effect of Time (Level Sequence) on performance ($\beta=0.090,SE=0.009,p<.001$). As illustrated in Figure 12, participants in all conditions demonstrated a steady improvement in weighted performance as the session progressed, indicating a continuous learning effect rather than fatigue. The positive slope suggests that users were able to handle increasingly difficult challenges (or maintain accuracy at higher difficulty levels) without a degradation in capacity.
This learning rate was consistent across all experimental conditions. The interaction between Time and Control Type was not significant ($p=.942$), nor was the three-way interaction between Time, Control Type, and Task Complexity ($p=.722$). This indicates that the rate of improvement was identical regardless of whether the difficulty was controlled by the model or the user, and regardless of the secondary task load.
While the model-controlled condition did not induce a steeper learning curve, it successfully supported the same rate of skill acquisition as in the self-controlled conditions. The absence of a negative interaction effect in the dual-task condition further confirms that the additional cognitive load did not inhibit the users’ ability to learn and improve at the primary task over time.
Figure 12: Linear regression of Weighted Performance trajectories over time. (a) Single Task condition (left) and (b) Dual Task condition (right). Red lines represent Model Control; Blue lines represent Self Control. The significant positive slope across all conditions ($p<.001$) indicates a consistent learning effect. Shaded regions represent 95% confidence intervals.
This continuous learning trajectory culminated in significant performance retention between the first and second exposures of each task. Repeated-measures ANOVA confirmed that Weighted Game Performance improved significantly in the second exposure for both Single-Task ($\Delta=+0.224,p<.001,d=0.87$) and Dual-Task ($\Delta=+0.146,p<.001,d=0.53$) conditions.
Raw Game Performance (unweighted) showed a more modest improvement, reaching significance only in the Dual-Task condition ($p=.016$). This discrepancy between weighted and raw metrics highlights the function of the adaptive system: as users learned (positive slope), the system increased the difficulty to match their new capacity. Consequently, while their raw score remained relatively stable, their difficulty-weighted performance increased because they were operating at a significantly higher difficulty level. This is confirmed by the analysis of Final Difficulty, which significantly increased (numerical value decreased) in the second Single-Task exposure ($p<.001,d=0.53$), indicating that the learning observed in the LMM was successfully converted into higher challenge acceptance.